OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 1) (5, 4)

OpenStudy (anonymous):

What you need to do first is find out the slope y2-y1 / x2 - x1 = 4-1 / 5 - 6 = 3/-1 = -3 y=-3x+b Are you with me so far?